Titan Football Manufacturing had the following operating results for 2016: sales = $19,930; cost of goods sold = $13,830; depreciation expense = $2,220; interest expense = $270; dividends paid = $700. At the beginning of the year, net fixed assets were $21,300, current assets were $3,090, and current liabilities were $1,920. At the end of the year, net fixed assets were $25,340, current assets were $3,580, and current liabilities were $2,010. The tax rate for 2016 was 30 percent.

a. What is net income for 2016? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Net income             $

b. What is the operating cash flow for 2016? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Operating cash flow             $

c. What is the cash flow from assets for 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

   

Cash flow from assets            $

Assume no new debt was issued during the year.

  

d. What is the cash flow to creditors for 2016? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Cash flow to creditors            $

e. What is the cash flow to stockholders for 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Cash flow to stockholders         
Answer #1

a)Net income=( sales-cost of goods sold-depreciaiton-interest)*(1-tax)
=(19930-13830-2220-270)*(1-30%)
=2527
b)operating cash flow=EBIT+depreciation-taxes
EBIT=sales-cost of goods sold-depreciaiton
=(19930-13830-2220)
=3880
Taxes=(EBIT-interest)*tax
=(3880-270)*30%
=1083
operating cash flow=3880+2220-1083=5017
c)Cash flow from assets = OCF – Change in Net working capital – Net capital spending
Change in Net working capital=(current asset end-current asset beg)-(current liab end-current liab beg)
=(3580-3090)-(2010-1920)
=400
Net capital spending= Net capital spend ending- net capital beg+depreciation
=25340-21300+2220
=6260
Cash flow from assets=5017-400-6260=-1643
d)Cash flow to creditors = Interest – new debt issued
=270-0
=270
e)Cash flow from assets = Cash flow to stockholders + Cash flow to creditors
Cash flow to stockholder=-1643-270=-1913